# Wetting on Silicone Surfaces

### v1

@@ III.1.2 Elastic
Wetting ridges also occur on elastomeric surfaces [150]. However, the mechanisms that govern the ”elastic wetting ridge” are more complex than those governing liquid ridges. Capillary forces around the perimeter of a wetting drop induce mechanical stress in the PDMS material $(\bm{\sigma})$. Consequently, the material around the three-phase contact line deforms to a wetting ridge, similar to the liquid wetting ridge. In addition to the surface tension of the PDMS surface, the elasticity of the network bulk contributes to the ridge geometry. Depending on the bulk stiffness, wetting ridges on elastomers can grow up to tens of microns. With the steady improvement of optical techniques, elastic wetting ridges have been resolved with an increasing level of detail. Such modern state-of-the-art techniques are e.g., shadowgraphy, interferometry, and laser scanning confocal microscopy (cf. Methods and Techniques, [53, 132, 151, 152], and Fig. 3.)
@@
$\bm{\sigma}\vec{n}-\Upsilon_{\mathrm{e}}\left(\nabla\cdot\vec{n}\right)\vec{n}=\vec{\gamma}_{w}.$ (6)
The spring-load system is a simple analogy to this wetting ridge where the resulting displacement - i.e., the height of the wetting ridge $\delta_{e,\mathrm{I}}$ [153, 154] - scales with the drop surface tension pulling on the surface per
$\delta_{e,\mathrm{I}}\propto\frac{\gamma_{w}}{\mu}\sin{\theta}.$ (7)
The compliance of the wetting ridge is characterized by $\mu^{-1}$, where $\mu$ is the shear modulus of PDMS. PDMS elastomers have a hydrophobic contact angle $\theta$ between $90^{\circ}-110^{\circ}$ [75, 76, 77, 78, 79, 80], Fig 1b. Here, the solid surface tension $\Upsilon_{\mathrm{e}}$ is considered symmetric around $x_{\mathrm{cl}}$. This, however, is not fulfilled as PDMS-air and PDMS-water interfaces have differing surface tension ($\Upsilon_{el}\approx 40~{}\mathrm{mN/m}$ and $\Upsilon_{e}\approx 20~{}\mathrm{mN/m}$). To match the asymmetric surface tension around the contact line, Eq. (6) can be adjusted with Heaviside functions.
Solutions of this system of equations (and variations) show that the competition between the bulk elasticity $(\mu)$ and the capillary point load $(\gamma_{w}\sin{\theta})$ produces shapes that are comparable with experimental measurements [132].
@@
$\delta_{e,\mathrm{II}}\propto\frac{\Upsilon_{e}}{\mu}.$ (8)
When the wetting drop induces an elastic wetting ridge where $\mathcal{O}(\delta_{e,\mathrm{I}})$ is below $\mathcal{O}(\delta_{e,\mathrm{II}})$ the Neumann configuration [Eq. (1)] is recovered. The ratio of the two elastic length scales $\left[\alpha=\delta_{e,\mathrm{I}}/\delta_{e,\mathrm{II}}\right]$ demarcates the two limits in which the stress can be treated:
$\alpha\propto\frac{\gamma_{w}}{\Upsilon_{e}}\sin{\theta}\begin{cases}\ll 1&\mathrm{Neumann}\text{-}\mathrm{limit,}\\ \gg 1&\mathrm{bulkelasticity}\text{-}\mathrm{limit}.\end{cases}$ (9)
@@
$\Upsilon_{\mathrm{e}}=\frac{\mathrm{d}}{\mathrm{d}\epsilon_{\Sigma}}\left[(1+\epsilon_{\Sigma})\gamma_{e}(\epsilon_{\Sigma})\right].$ (10)
Eq. (10) is known as the Shuttleworth equation [157, 158]. While originally introduced in 1950, it obtained recent momentum in the wetting community. The Shuttleworth effect was utilized to show anisotropic wetting behavior on unidirectionally stretched PDMS surfaces [159, 160, 161] and differing dependencies of $\Upsilon_{\mathrm{e}}$ and $\gamma_{e}$ on the cross-link density [162]. Other studies found contradictory behavior and the absence of a $\gamma_{e}(\epsilon_{\Sigma})$ dependency [163]. Their observations were rationalized by only partially stretched elastomers (i.e., below the max. chain length) or oil-swollen elastomers where mobile chains at the surface enable the relaxation of strain-induced local molecular surface inhomogeneities [164, 165, 113]. In that sense, the PDMS elastomer surfaces resemble rather a liquid interface than a solid one. However, surface modifications, e.g. oxidation of surface molecules, can create a thin ”surface skin” that is chemically and mechanically different from the bulk PDMS [166, 167]. Upon stretching, the oxidized skin breaks, and non-oxidized PDMS molecules replenish the gap. Such surface-modified PDMS elastomers exhibit a strain-dependent surface energy/tension per Eq. (10).
@@ III.1.3 Molecular
A wetting ridge on a SOCAL surface is tiny but possible: chains may stretch around the contact line, or ungrafted chains migrate freely to the contact line. Because of the nanometric thickness of the SOCAL surface, direct experimental observations of ”molecular wetting ridges” have not been demonstrated. However, theoretical calculations - both on a molecular level and in a mean-field framework - predict their formation [169, 170, 171, 133].
The molecular wetting ridge is, again, a result of the interplay of the capillary forces induced by the wetting drop and the material response of the surface. However, the motivation of the linear elasticity framework from a continuum mechanical point of view is less obvious as the length scales are much smaller. Deformations occur (at most) over the length of a chain (collapsed to stretched) in contrast to elastomers that undergo deformations far exceeding the lengthscales of the spacing between crosslinking nodes. Further, the chain configuration introduces isotropy as the chains are not laterally reticulated. Therefore, the ridge deformation on SOCAL surfaces is usually considered in a thermodynamic framework of the individual polymer chains. The restoring force on the molecular chain level originates from entropy [59]. Stretched chains have a reduced conformational state. Thus, the entropy is low. This state is undesirable for the grafted chain. It strives to increase the possible conformations and maximize its entropy by recovering the unstretched state [172, 173]. Notably, the thermodynamic consideration for a swollen chain recovers a linear relation between the deformation and the restoring force per chain
@@
$\delta_{m}\propto\frac{\gamma_{w}nb^{2}}{\sqrt{\sigma}k_{B}T}\sin{\theta},$ (12)
where $\sigma$ is the grafting density for the SOCAL surface ($\sigma\approx 0.1~{}\text{chains/nm}^{2}$ [121]).
III.2 Free PDMS Chains and Cloaking
@@ III.2.2 Cloaking Layer
The surface tension of a wetting drop yields the formation of a wetting ridge and the accumulation of free chains (PDMS oil). A sharp three-phase contact line at the tip of the ridge implies a (Neumann) balance [Eq. (1)] of the competing surface tensions of the water ($\gamma_{w}$), the oil ($\gamma_{o}$), and water/oil ($\gamma_{ow}$). The Neumann balance is fulfilled with a negative ”spreading coefficient”:
@@
Generally, $S\leq 0$ indicates the presence of a contact line (partial wetting), and $S>0$ expresses the absence of a contact line in equilibrium (complete wetting). As all three interfaces are liquid, their surface tensions are easily accessible with direct measurements (cf. Shadowgraphy in Methods and Techniques), and their values are listed in Tab. 1. Notably, for these values, the spreading coefficient is positive $(7-16~{}\mathrm{mN/m})$, and thus, PDMS oil tends to spread on top of the water drop [48]. The PDMS engulfs the drop with a thin layer, Fig. 4a,b. This behavior is called ”cloaking”, and significantly changes the wetting configuration [181]. Due to the cloak, the surface energy of the drop cap changes, affecting the total free energy and consequently the wetting affinity of the drop and the contact angle, respectively. A cloaking layer of PDMS oil makes the Neumann configuration impossible. Instead, the effective surface tension of the drop changes.
@@
water-air $\gamma_{w}$ $72$ (self-measured)
@@
Naively, the ”cloaked” surface tension $\gamma_{c}$ of the drop effectively sums up from two interfaces (water-oil and oil-air) that replaces the uncloaked surface (water-air), i.e., $\gamma_{c}=\gamma_{ow}+\gamma_{o}$. However, the concepts of surface tension and surface energy consider shared surfaces of two bulk phases. Surface tension and energy result from excess energy that decays at some length scale towards the bulk phases. The length (or the vicinity) in which the excess energy decays is also reflected in the stress isotropy, being isotropic in the bulk phases, and anisotropic near the interface. Following this reasoning, $\gamma_{c}$ is only the sum of $\gamma_{ow}$ and $\gamma_{o}$ when the cloaking layer is thick enough and the surface tension has ”bulk properties”. However, the layer can be substantially thinner, especially upon cloak formation. In this case, the two interfaces (water-oil and oil-ambient) of the thin cloak layer start to interact with each other via the disjoining pressure $\Pi(h_{c})$ yielding stabilizing stress contributions which depend on the thickness of the cloak $h_{c}$. The stabilizing disjoining pressure contribution adds to the surface tension of the cloak with [185, 186, 187]
@@
Figure 4: Sessile drop on lubricated surface, cloaked by PDMS oil. a) Three-dimensional view of PDMS oil on the drop, obtained by laser scanning confocal microscopy. b) Side view of the angular averaged 3D stack. Images adapted from [133] c) Temporal evolution of pendant drop on PDMS elastomer with swollen with $200~{}\mathrm{mPa~{}s}$ (red) and $350~{}\mathrm{mPa~{}s}$ (green) PDMS oil. The shaded region shows standard deviation of experimental repeats. Data adapted from [189].
@@ III.3.1 Liquid
Sliding drops experience friction, stemming from shear dissipation. On rigid surfaces, the source of dissipation is either found at the three-phase contact line (for small speeds) or at the footprint of the drop where a hydrodynamic boundary layer forms (at higher speeds). On PDMS surfaces, the main source of dissipation, however, can be localized in the PDMS layer instead of the drop. This is the case when the lubricant is more viscous than the drop [190]. The wetting ridge at the three-phase contact line moves with the drop and the accumulated material in the ridge is constantly reorganized, Fig. 5a. During reorganization, the material undergoes shear and strain that dissipates energy due to the viscous/viscoelastic nature of the PDMS material. Note, that shear is not only evoked in the ridge but on all parts of the PDMS surface that are in contact with the drop (i.e., the drop footprint); however, the dissipation in the wetting ridge is the dominant source of sliding friction. Hence, to quantify the sliding friction, it is sufficient to only focus on the wetting ridge.
@@
In general, the dissipation $P_{\mathrm{diss}}=\int_{V_{\mathrm{ridge}}}\bm{\sigma}:\dot{\bm{\epsilon}}dV$, depends on the volume of the wetting ridge $V_{\mathrm{ridge}}$, the strain rate $\dot{\bm{\epsilon}}$ (i.e., the rate of reorganization), and the material stress $\bm{\sigma}$. The stress in the wetting ridge follows from material laws, and thus, changes with the fabric of the different surface types. PDMS oils are Newtonian fluids that show a linear relationship between stress and shear rate, $\bm{\sigma}=\eta\dot{\bm{\epsilon}}$. If the oil is very viscous, the drop mobility is significantly hampered [190, 191]. Still, lubrication enables higher drop mobility on lubricated compared to non-lubricated surfaces which tend to be rougher and have a much larger number of surface defects that slow down the drop [24]. The shear stress in the liquid wetting ridge scales as $\sigma\propto\eta v/h_{r}$. The dissipation rate is found by integrating the viscous stress over the ridge area. Furthermore, we must account for the variation of dynamic contact angle of the wetting ridge with speed, Fig 5b,c), which leads to, $P_{\mathrm{diss}}\propto v\beta\phi r\left(\eta v/\theta_{o}\right)$. $\phi$ accounts for the texture density which is important as only texture contacts along the perimeter add to the overall friction. On parts where the ridge contacts oil, it is assumed that no significant shear builds up and can, therefore, be neglected. $\beta=\ln{(h_{r}/\epsilon)}$ and $\epsilon$ account for cut-off length scale (taken to be $\approx 100~{}\mathrm{nm}$) [192]. $\theta_{o}$ is the opening angle, formed between the horizontal plane and the wetting ridge. This angle changes with the speed according to Tanner’s law $\theta_{o}\approx\left(\beta\eta v/\gamma\right)^{1/3}$ [193]. Including Tanner’s law into the expression for the dissipative power yields $P_{\mathrm{diss}}\propto v\phi r\left(\beta\eta v/\gamma\right)^{2/3}$. The ratio between the viscous stress and the surface force in the liquid wetting ridge are condensed in the capillary number $\mathrm{Ca}=\eta v/\gamma$. Hence, the dissipative scaling can be written compactly as
@@
On flat lubricated surfaces, the ridge meniscus behaves analogously to a Landau Levich Derjaguin film: the viscous stress elevates the film in the vicinity $l_{\mathrm{dyn}}$ behind the advancing and the receding wetting ridge. The elevated height scales as $h_{\mathrm{dyn}}\propto r\mathrm{Ca}^{2/3}$, and the vicinity length as $l_{\mathrm{dyn}}\propto r\mathrm{Ca}^{1/3}.$ Depending on whether considering the advancing or receding ridge, $\mathrm{Ca}$ is formed with $\gamma_{ow}$ and $\gamma_{o}$, respectively. As many oil surfaces utilize a porous structure to maintain better retention of the oil film, $h_{\mathrm{dyn}}$ only forms when this height exceeds the height of the porous structure $h_{\mathrm{p}}$ [90]. Combining the dynamic $h_{\mathrm{dyn}}$ scaling with this inequality, produces the critical capillary number $\mathrm{Ca}^{*}=(h_{\mathrm{p}}/r)^{3/2}$ that needs to be exceeded to form $h_{\mathrm{dyn}}$ [191]. The associated (secondary) dissipation rate within this domain scales (non-linearly in $v$) as
@@
Figure 5: Dynamic geometries of liquid wetting ridge during drop sliding. The oil film underneath the drop and behind the drop, both, resemble a Landau-Levich film. The dynamic film thickness a) underneath the drop front $h_{\mathrm{dyn,f}}$, and b) behind the drop rear end $h_{\mathrm{dyn,b}}$, both display the scaling $\propto\mathrm{Ca}^{2/3}$, characteristic for Landau-Levich films. Data in b) and c) adapted from [50] and insets in b) adapted from [190].
Since the scaling between power and the capillary number is the same for both dissipation in the Landau Levich film and at the foot of the wetting ridge, most experiments have the same scaling regardless of the dissipation mechanism. While this makes the scaling law remarkably universal, it also makes it difficult to directly deduce which of the two mechanisms is dominant.
At high capillary numbers (Ca $\gtrsim 10^{-2}$), a different scaling law was observed between friction on velocity, with a lower exponent on the velocity. In this regime, drops begin to move more rapidly. The mechanism for this regime is still unknown. A tentative explanation that has been proposed is that the wetting ridge shape does not have time to fully develop in this regime [194].
Figure 6: Dissipation-speed profile ($P_{\mathrm{diss}}/r\propto v^{k}$) on a) liquid-infused silicone oil (green) $\eta=100~{}\mathrm{mPa~{}s}$ and perfluorinated oil (red) $\eta=30~{}\mathrm{mPa~{}s}$ ($k\approx 1.66$), b) $\mu=250~{}\mathrm{kPa}$ elastomer ($k\approx 1.23$), and c) $6~{}\mathrm{kDa}$ chain SOCAL surfaces ($k\approx 1.25$). a) Bright colors correspond to cases where $\mathrm{Ca}^{*}<\mathrm{Ca}$ and dark colors to cases where $\mathrm{Ca}^{*}>\mathrm{Ca}$, with $\mathrm{Ca}^{*}=(h_{\mathrm{p}}/r)^{3/2}$.b) The colors correspond to various coating thickness: green - $1100~{}\mathrm{nm}$, black berry - $350~{}\mathrm{nm}$, red - $90~{}\mathrm{nm}$, yellow - $55~{}\mathrm{nm}$. c) Data points were measured on SOCAL surface, with chains: pink - grafted from, blue - grafted to substrate (c.f. PDMS based surfaces - structure and properties). Data reproduced from a) red [195] green [190, 191], b) [196], and c) green [197] rest [64].

### v2

@@ III.1.2 Elastic
Wetting ridges also occur on elastomeric surfaces [176]. However, the mechanisms that govern the “elastic wetting ridge” are more complex than those governing liquid ridges. Capillary forces around the perimeter of a wetting drop induce mechanical stress in the PDMS material $(\mathbf{\sigma})$. Consequently, the material around the three-phase contact line deforms to a wetting ridge, similar to the liquid wetting ridge. In addition to the surface tension of the PDMS surface, the elasticity of the network bulk contributes to the ridge geometry. Depending on the bulk stiffness, wetting ridges on elastomers can grow up to tens of microns. With the steady improvement of optical techniques, elastic wetting ridges have been resolved with an increasing level of detail. Such modern state-of-the-art techniques are e.g., shadowgraphy, interferometry, and laser scanning confocal microscopy (cf. Methods and Techniques, [55, 177, 178, 68], and Fig. 3.)
@@
Figure 3: Different manifestations of the wetting ridge on different PDMS surfaces. a) Liquid wetting ridge on a liquid-infused surface. Neumann balance at the three-phase contact line (box) and the apparent contact angle $\theta_{app}$. The right half illustrates the starved limit and the pseudo-Young’s balance. Macroscopically, the origin of the surface tension vectors (point $A$) aligns with the substrate baseline. The image on the right shows a liquid wetting ridge, taken with a fluorescence confocal microscope, adapted from [118]. b) Elastic wetting ridge on an elastomeric surface. Bulk elasticity creates a compliance force $f_{el}\propto G\lambda_{e,\mathrm{I}}$, resulting in wetting ridges smaller than the liquid ones. The right half shows material dimpling around the wetting ridge when the surface coating is very thin. The right image shows an elastic wetting ridge, taken with an x-ray microscope on elastomeric PDMS ($3~{}\mathrm{kPa}$), adapted from [177]. c) Molecular wetting ridge on a SOCAL surface. The entropic force of the PDMS chain $f_{m}$ limits the surface tension-induced chain stretch. Right half shows surface swelling when liquid-PDMS affinity/solubility is high. The right image shows a molecular wetting ridge, computed with coarse-grained molecular dynamics simulation, adapted from [88].
@@
$\mathbf{\sigma}\vec{n}-\Upsilon_{e}\left(\nabla\cdot\vec{n}\right)\vec{n}=\vec{\gamma}_{w}.$ (6)
The spring-load system is a simple analogy to this wetting ridge where the resulting displacement - i.e., the height of the wetting ridge $\lambda_{e,\mathrm{I}}$ [179, 180] - scales with the drop surface tension pulling on the surface per
$\lambda_{e,\mathrm{I}}\propto\frac{\gamma_{w}}{G}\sin{\theta}.$ (7)
The compliance of the wetting ridge is characterized by $G^{-1}$, where $G$ is the shear modulus of PDMS. PDMS elastomers have a hydrophobic contact angle $\theta$ between $90^{\circ}-110^{\circ}$ [101, 102, 103, 104, 105, 106], Fig 1b. Here, the solid surface tension $\Upsilon_{e}$ is considered symmetric around $x_{\mathrm{cl}}$. This, however, is not fulfilled as PDMS-air and PDMS-water interfaces have differing surface tension ($\Upsilon_{el}\approx 40~{}\mathrm{mN/m}$ and $\Upsilon_{e}\approx 20~{}\mathrm{mN/m}$). To match the asymmetric surface tension around the contact line, Eq. (6) can be adjusted with Heaviside functions. Note, the asymmetric surface tensions (dry vs. wetted) is not only invoked by water but is general for most liquids.
Solutions of this system of equations (and variations) show that the competition between the bulk elasticity $(G)$ and the capillary point load $(\gamma_{w}\sin{\theta})$ produces shapes that are comparable with experimental measurements [177].
@@
$\lambda_{e,\mathrm{II}}\propto\frac{\Upsilon_{e}}{G}.$ (8)
When the wetting drop induces an elastic wetting ridge where $\mathcal{O}(\lambda_{e,\mathrm{I}})$ is below $\mathcal{O}(\lambda_{e,\mathrm{II}})$ the Neumann configuration [Eq. (1)] is recovered. The ratio of the two elastic length scales the strain $\left[\epsilon=\lambda_{e,\mathrm{I}}/\lambda_{e,\mathrm{II}}\right]$ and demarcates the two limits in which the stress can be treated:
$\epsilon\propto\frac{\gamma_{w}}{\Upsilon_{e}}\sin{\theta}\begin{cases}\ll 1&\mathrm{Neumann}\text{-}\mathrm{limit,}\\ \gg 1&\mathrm{bulkelasticity}\text{-}\mathrm{limit}.\end{cases}$ (9)
@@
$\Upsilon_{e}=\frac{\mathrm{d}}{\mathrm{d}\epsilon_{\Sigma}}\left[(1+\epsilon_{\Sigma})\gamma_{e}(\epsilon_{\Sigma})\right].$ (10)
Eq. (10) is known as the Shuttleworth equation [183, 184]. While originally introduced in 1950, it obtained recent momentum in the wetting community. The Shuttleworth effect was utilized to show anisotropic wetting behavior on unidirectionally stretched PDMS surfaces [185, 186, 187] and differing dependencies of $\Upsilon_{e}$ and $\gamma_{e}$ on the cross-link density [188]. Other studies found contradictory behavior and the absence of a $\gamma_{e}(\epsilon_{\Sigma})$ dependency [189]. Their observations were rationalized by only partially stretched elastomers (i.e., below the max. chain length) or oil-swollen elastomers where mobile chains at the surface enable the relaxation of strain-induced local molecular surface inhomogeneities [190, 191, 78]. In that sense, the PDMS elastomer surfaces resemble rather a liquid interface than a solid one. However, surface modifications, e.g. oxidation of surface molecules, can create a thin “surface skin” that is chemically and mechanically different from the bulk PDMS [192, 193]. Upon stretching, the oxidized skin breaks, and non-oxidized PDMS molecules replenish the gap. Such surface-modified PDMS elastomers exhibit a strain-dependent surface energy/tension per Eq. (10).
@@ III.1.3 Molecular
A wetting ridge on a SOCAL surface is tiny but possible: chains may stretch around the contact line, or ungrafted chains migrate freely to the contact line. Because of the nanometric thickness of the SOCAL surface, direct experimental observations of “molecular wetting ridges” have not been demonstrated. However, theoretical calculations - both on a molecular level and in a mean-field framework - predict their formation [160, 161, 195, 88].
The molecular wetting ridge is, again, a result of the interplay of the capillary forces induced by the wetting drop and the material response of the surface. However, the motivation of the linear elasticity framework from a continuum mechanical point of view is less obvious as the length scales are much smaller. Since water is a poor solvent for PDMS [97, 98, 99], the grafted chains are in a collapsed state, underneath the drop, Fig. 3c left box. Deformations occur (at most) over the length of a chain (collapsed to stretched) in contrast to elastomers that undergo deformations far exceeding the lengthscales of the spacing between crosslinking nodes. Organic solvents are much better solvents[40] and enable chain stretching, leading to smaller wetting ridges, Fig. 3 center box. Further, the chain configuration introduces isotropy as the chains are not laterally reticulated. Therefore, the ridge deformation on SOCAL surfaces is usually considered in a thermodynamic framework of the individual polymer chains. The restoring force on the molecular chain level originates from entropy [61]. Stretched chains have a reduced conformational state. Thus, the entropy is low. This state is undesirable for the grafted chain. It strives to increase the possible conformations and maximize its entropy by recovering the unstretched state [196, 197]. Notably, the thermodynamic consideration for a swollen chain recovers a linear relation between the deformation and the restoring force per chain
@@
$\lambda_{m}\propto\frac{\gamma_{w}nb^{2}}{\sqrt{\rho}k_{B}T}\sin{\theta},$ (12)
where $\rho$ is the grafting density for the SOCAL surface $\left[\rho\approx 0.1~{}\text{chains/nm}^{2}\right]$ [143].
III.2 Free Chains, Cloaking, and Swelling
@@ III.2.2 Cloaking Layer
The surface tension of a wetting drop yields the formation of a wetting ridge and the accumulation of free chains (PDMS oil) [79, 81]. A sharp three-phase contact line at the tip of the ridge implies a (Neumann) balance [Eq. (1)] of the competing surface tensions of the water ($\gamma_{w}$), the oil ($\gamma_{o}$), and water/oil ($\gamma_{ow}$). The Neumann balance is fulfilled with a negative “spreading coefficient”:
@@
Generally, $S\leq 0$ indicates the presence of a contact line (partial wetting), and $S>0$ expresses the absence of a contact line in equilibrium (complete wetting). As all three interfaces are liquid, their surface tensions are easily accessible with direct measurements (cf. Shadowgraphy in Methods and Techniques) and for water drops, values are listed in Tab. 1. Notably, for these values, the spreading coefficient is positive $(7-16~{}\mathrm{mN/m})$, and thus, PDMS oil tends to spread on top of the water drop [50]. The PDMS engulfs the drop with a thin layer, Fig. 4a,b. This behavior is called “cloaking”, and significantly changes the wetting configuration [89]. Due to the cloak, the surface energy of the drop cap changes, affecting the total free energy and consequently the wetting affinity of the drop and the contact angle, respectively. A cloaking layer of PDMS oil makes the Neumann configuration impossible. Instead, the effective surface tension of the drop changes.
@@
water-air $\gamma_{w}$ $72$ [90]
@@
Naively, the “cloaked” surface tension $\gamma_{c}$ of the drop effectively sums up from two interfaces (water-oil and oil-air) that replaces the uncloaked surface (water-air), i.e., $\gamma_{c}=\gamma_{ow}+\gamma_{o}$. However, the concepts of surface tension and surface energy consider shared surfaces of two bulk phases. Surface tension and energy result from excess energy that decays at some length scale towards the bulk phases [207]. The length (or the vicinity) in which the excess energy decays is also reflected in the stress isotropy, being isotropic in the bulk phases, and anisotropic near the interface [208]. Following this reasoning, $\gamma_{c}$ is only the sum of $\gamma_{ow}$ and $\gamma_{o}$ when the cloaking layer is thick enough and the surface tension has “bulk properties”. However, the layer can be substantially thinner, especially upon cloak formation [52]. In this case, the two interfaces (water-oil and oil-ambient) of the thin cloak layer start to interact with each other via the disjoining pressure $\Pi(h_{c})$ yielding stabilizing stress contributions which depend on the thickness of the cloak $h_{c}$. The stabilizing disjoining pressure contribution adds to the surface tension of the cloak with [209, 210, 211]
@@
Figure 4: Sessile drop on lubricated surface, cloaked by PDMS oil. a) Three-dimensional view of PDMS oil on the drop, obtained by laser scanning confocal microscopy. b) Side view of the angular averaged 3D stack. Images adapted from [88] c) Temporal evolution of pendant drop on PDMS elastomer with swollen with $200~{}\mathrm{mPa~{}s}$ (red) and $350~{}\mathrm{mPa~{}s}$ (green) PDMS oil. The shaded region shows the standard deviation of experimental repeats. Data adapted from [213].
III.2.3 Swelling
Water and other polar liquids typically have a poor affinity with PDMS (cf. Structure and Properties). Other liquids (in particular organic ones) however, interact much stronger with PDMS. When such organic liquids wet a PDMS surface (LIS, elastic, or SOCAL), the wetting drop can ”swell” the underlying PDMS substrate. The extent of swelling depends on the affinity between the wetting liquid and the PDMS. For some liquids (e.g. toluene or hexane) the liquid fully immerses into the PDMS, increasing the volume by more than 220 % [40]. Swelling alters the material properties of the PDMS surface, e.g. the surface tension [214, 49, 43], the shear modulus [215], the viscosity, etc. Hence, swelling directly influences the static and dynamic wetting behavior, i.e., the contact angle $\theta$, the hysteresis $\Delta\theta$, and the dissipation (cf. Dynamic Wetting Ridge).
When the wetting drop moves, the swelling kinetics can compete with the sliding of the drop. This leads to a steady state where the degree of swelling is not equilibrated but depends on the sliding speed of the drop. In this state, the surface underneath the front and rear of the drop swell to different degrees, and the PDMS wettability exhibits local gradients. This induces a contact angle hysteresis that does not depend on the surface inhomogeneity but the swelling kinetics, the drop size, and the sliding speed [43].
@@ III.3.1 Liquid
Sliding drops experience friction, stemming from shear dissipation. On rigid surfaces, the source of dissipation is either found at the three-phase contact line (for small speeds) or at the footprint of the drop where a hydrodynamic boundary layer forms (at higher speeds). On PDMS surfaces, the main source of dissipation, however, can be localized in the PDMS layer instead of the drop. This is the case when the lubricant is more viscous than the drop [74]. The wetting ridge at the three-phase contact line moves with the drop and the accumulated material in the ridge is constantly reorganized, Fig. 5a. During reorganization, the material undergoes shear and strain that dissipates energy due to the viscous/viscoelastic nature of the PDMS material. Note, that shear is not only evoked in the ridge but on all parts of the PDMS surface that are in contact with the drop (i.e., the drop footprint); however, the dissipation in the wetting ridge is the dominant source of sliding friction [83, 74, 84, 85]. Hence, to quantify the sliding friction, it is sufficient to only focus on the wetting ridge.
@@
In general, the dissipation $P_{\mathrm{diss}}=\int_{V_{\mathrm{ridge}}}\mathbf{\sigma}:\dot{\mathbf{\epsilon}}dV$, depends on the volume of the wetting ridge $V_{\mathrm{ridge}}$, the strain rate $\dot{\mathbf{\epsilon}}$ (i.e., the rate of reorganization), and the material stress $\mathbf{\sigma}$. The stress in the wetting ridge follows from material laws, and thus, changes with the fabric of the different surface types. PDMS oils with $\eta<20~{}\mathrm{Pa~{}s}$ are Newtonian fluids that show a linear relationship between stress and shear rate, $\mathbf{\sigma}=\eta\dot{\mathbf{\epsilon}}$. If the oil is very viscous it can exhibit shear-thinning behavior [216], however, the drop mobility is significantly hampered in the Newtonian low-speed regime[74, 84]. Still, lubrication enables higher drop mobility on lubricated compared to non-lubricated surfaces which tend to be rougher and have a much larger number of surface defects that slow down the drop [26]. The shear stress in the liquid wetting ridge scales as $\sigma\propto\eta v/h_{r}$. The dissipation rate is found by integrating the viscous stress over the ridge area. Furthermore, we must account for the variation of dynamic contact angle of the wetting ridge with speed, Fig 5b,c), which leads to, $P_{\mathrm{diss}}\propto v\beta\phi r\left(\eta v/\theta_{o}\right)$. $\phi$ accounts for the texture density which is important as only texture contacts along the perimeter add to the overall friction. On parts where the ridge contacts oil, it is assumed that no significant shear builds up and can, therefore, be neglected. $\beta=\ln{(h_{r}/\alpha)}$ and $\alpha\approx 100~{}\mathrm{nm}$ account for a cut-off length [217]. $\theta_{o}$ is the opening angle, formed between the horizontal plane and the wetting ridge. This angle changes with the speed according to Tanner’s law $\theta_{o}\approx\left(\beta\eta v/\gamma\right)^{1/3}$ [218]. Including Tanner’s law into the expression for the dissipative power yields $P_{\mathrm{diss}}\propto v\phi r\left(\beta\eta v/\gamma\right)^{2/3}$. The ratio between the viscous stress and the surface force in the liquid wetting ridge are condensed in the capillary number $\mathrm{Ca}=\eta v/\gamma$. Hence, the dissipative scaling can be written compactly as
@@
On flat lubricated surfaces, the ridge meniscus behaves analogously to a Landau Levich Derjaguin film: the viscous stress elevates the film in the vicinity $l_{\mathrm{dyn}}$ behind the advancing and the receding wetting ridge [219, 83]. The elevated height scales as $h_{\mathrm{dyn}}\propto r\mathrm{Ca}^{2/3}$, and the vicinity length as $l_{\mathrm{dyn}}\propto r\mathrm{Ca}^{1/3}.$ Depending on whether considering the advancing or receding ridge, $\mathrm{Ca}$ is formed with $\gamma_{ow}$ and $\gamma_{o}$, respectively. As many oil surfaces utilize a porous structure to maintain better retention of the oil film, $h_{\mathrm{dyn}}$ only forms when this height exceeds the height of the porous structure $h_{\mathrm{p}}$ [116]. Combining the dynamic $h_{\mathrm{dyn}}$ scaling with this inequality, produces the critical capillary number $\mathrm{Ca}^{*}=(h_{\mathrm{p}}/r)^{3/2}$ that needs to be exceeded to form $h_{\mathrm{dyn}}$ [84]. The associated (secondary) dissipation rate within this domain scales (non-linearly in $v$, Fig. 6a) as
@@
Figure 5: Dynamic geometries of liquid wetting ridge during drop sliding. The oil film underneath the drop and behind the drop, both, resemble a Landau Levich Derjaguin film. The dynamic film thickness a) underneath the drop front $h_{\mathrm{dyn,f}}$, and b) behind the drop rear end $h_{\mathrm{dyn,b}}$, both display the scaling $\propto\mathrm{Ca}^{2/3}$, characteristic for Landau Levich Derjaguin films. Data in b) and c) adapted from [52] and insets in b) adapted from [74].
Since the scaling between power and the capillary number is the same for both dissipation in the Landau Levich Derjaguin film and at the foot of the wetting ridge, most experiments have the same scaling regardless of the dissipation mechanism. While this makes the scaling law remarkably universal, it also makes it difficult to directly deduce which of the two mechanisms is dominant.
At high capillary numbers ($\mathrm{Ca}>\approx 10^{-2}$), a different scaling law was observed between friction on velocity, with a lower exponent on the velocity. In this regime, drops begin to move more rapidly. The mechanism for this regime is still unknown. A tentative explanation that has been proposed is that the wetting ridge shape does not have time to fully develop in this regime [220].
Figure 6: Dissipation-speed profile ($P_{\mathrm{diss}}/r\propto v^{k}$) on a) liquid-infused silicone oil (green) $\eta=100~{}\mathrm{mPa~{}s}$ and perfluorinated oil (red) $\eta=30~{}\mathrm{mPa~{}s}$ ($k\approx 1.66$), b) $G=250~{}\mathrm{kPa}$ elastomer ($k\approx 1.23$), and c) $6~{}\mathrm{kDa}$ chain SOCAL surfaces ($k\approx 1.25$). a) Bright colors correspond to cases where $\mathrm{Ca}^{*}<\mathrm{Ca}$ and dark colors to cases where $\mathrm{Ca}^{*}>\mathrm{Ca}$, with $\mathrm{Ca}^{*}=(h_{\mathrm{p}}/r)^{3/2}$.b) The colors correspond to various coating thicknesses: green - $1100~{}\mathrm{nm}$, blackberry - $350~{}\mathrm{nm}$, red - $90~{}\mathrm{nm}$, yellow - $55~{}\mathrm{nm}$. c) Data points were measured on SOCAL surface, with chains: pink - grafted from, blue - grafted to substrate (c.f. PDMS based surfaces - structure and properties). Data reproduced from a) red [83] green [74, 84], b) [76], and c) green [221] rest [222].